On November 15 of the current year, the account balances in Orlole Partnership were Cash $8.650 Land $19.000D Dupuis Capitalpls show all calculation tx
Answer #1

Gain on realization $3,600 will be shared among all partners in the profit sharing ratio i.e. 1:1:1

Following entry is needed to transfer Gain on realization to partners

Journal

Date Account Debit Credit
Nov. 15 Gain on realization 3,600
B. Veitch, capital 1,200
V. Dueck, capital 1,200
D. Dupuls, capital 1,200

Cash to be paid to B. Veitch = B. Veitch, capital + Share of Gain on realization

= 4,450 + 1,200

= $5,650

Cash to be paid to V. Dueck = V. Dueck, capital + Share of Gain on realization

= 10,400 + 1,200

= $11,600

Cash to be paid to D. Dupuls = D. Dupuls, capital + Share of Gain on realization

= 12,800 + 1,200

= $14,000

Final cash balance = Beginning cash + Cash received from sale of Land

= 8,650 + 22,600

= $31,250

Following journal entry will be made for final cash distribution to partners:

Journal

Date Account Debit Credit
Nov. 15 B. Veitch, capital 5,650
V. Dueck, capital 11,600
D. Dupuls, capital 14,000
Cash 31,250

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.